### v2

# Emancipatory Information Retrieval

Abstract
Our world today is facing a confluence of several mutually reinforcing crises each of which intersects with concerns of social justice and emancipation. This paper is a provocation for the role of computer-mediated information access in our emancipatory struggles. We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term "emancipatory" here signifies the moral concerns of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. To develop an emancipatory research agenda for IR, in this paper we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. We challenge the field of information retrieval (IR) research to embrace humanistic values and commit to universal emancipation and social justice as part of our research.
1 Introduction
“The old world is dying, and the new world struggles to be born: now is the time of monsters.”
– Antonio Gramsci
Our world in crises.
Our world today is facing a confluence of several mutually reinforcing crises pushing us universally, but notably not uniformly, towards precarity. The recent COVID-19 pandemic did not only result in millions of deaths [87] and disrupt everyday lives globally; it also laid bare and further reinscribed existing political, economic, and social hierarchies that are saliently colonial and racialized [100], gendered [28, 84], casteist [94], and ableist [83]. Likewise, the dangers of recent and impending climate catastrophes as well as the discourse around mitigative actions are inextricably intertwined with concerns of racial [117, 136, 102, 116], gender [52, 1], and disability [37] justice. Increasing global wealth and income inequality [30] both mirrors historical dimensions of marginalization and is contributing to erosion of democracies worldwide [106]. Globally, we are still reckoning with the ongoing legacies of settler-colonialism [139] and present-day neocolonialism [98], recontesting women’s rights to bodily autonomy [17], and witnessing growing assaults on transgender people’s right to simply exist [49, 66].
The range of crises we face in the twenty-first century requires us to reckon with multiple intersecting and co-constituting processes of hierarchical organization of our society and as a consequence the systemic dehumanization, marginalization, and oppression of those designated to its bottom rung. Therefore, any attempt at addressing the crises of our times must begin with recommitment to the principles of social justice, universal humanization of all peoples, and emancipatory struggles against social, economic, and political oppression. This paper is a provocation for the role of computer-mediated information access in these struggles. Our aim is to challenge the field of information retrieval research to reimagine itself as what it can (or, ought to) be if grounded in humanistic values and committed to universal emancipation and social justice, and to encourage the community to explicitly articulate and make conscious choices about the values and incentives that shapes our field of computing.
The information reaches out.
Information retrieval (IR), the field of computing that concerns with the design and study of information access systems, such as web search and recommender systems, have their roots in library science and linguistics [63]. Early IR research was primarily concerned with the design of effective indexing mechanisms that enable fast pattern-based lookup of text from large corpora. The modern-day successors of these systems, however, go far beyond serving the function of simple lookup. Modern search systems observe and model collective behaviors of how their users interact with retrieved information to infer intent behind user queries, and to predict users’ future needs and actions. Some of these systems may proactively surface new information and breaking news, recommend commercial products and services, and even suggest activities and actions for the user to undertake. Conversational systems, such ChatGPT, [^1] https://chat.openai.com/ Microsoft Copilot, [^2] https://copilot.microsoft.com/ and Google Gemini [^3] https://gemini.google.com/, may be entrusted to summarize retrieved information in response to users’ search queries, effectively shifting the burden of assessing the relevance and trustworthiness of retrieved information and their contextual interpretation from the user to the statistical models underlying these systems. These systems affect what information and perspectives receive exposure which shapes consumer behavior, political discourse, and culture [55, 54, 58], and directly influence material outcomes for people [113, 39].
Today’s search and recommender systems are, therefore, much more than passive lookup tools. These systems have become embedded in our online social discourse, both shaping and being shaped by our collective values, cultural identities, sociopolitical beliefs, and shared understanding of our environment and ourselves. We could choose to see this transformation of information access systems from passive lookup tools to agents of influence as the information reaching out and permeating every aspect of our individual and social lives—akin to how Grudin [56] described the transition of computers from applications in narrow specialist domains to their present-day ubiquitous presence in our lives as “the computer reaching out”.
Importantly, these systems and platforms may privilege certain perspectives and ways of knowing over others that entrench social hierarchies and intensify oppression. Or, alternatively we could imagine future systems to serve to raise up marginalized voices and challenge the status quo. Therefore, they represent sites of conflict between exploitation and liberation. When the stakes are so high for our collective futures and wellbeing, how should we then reflect on and reconcile the implications of IR research and development on our communities and our collective liberation? In this paper, we assert that the field of IR would benefit from critically reflecting on how our work relates to current sociopolitical orders and our possible sociotechnical futures, and argues for reimagining IR research grounded explicitly in the adoption of humanistic and emancipatory values and aspirations.
2 What is emancipatory IR?
We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term “emancipatory” here signifies the moral concerns—central to this field of study—of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. It recognizes information and access to information both as profoundly political, and enacts research and development activities—such as, formalization, theorization, design, experimentation, publishing, deployment, open sourcing, and community building—in service of universal struggles against all forms of structural oppression. It rejects the techno-deterministic premise that there is a single pre-determined path forward for information access technology development, and challenges its practitioners to employ reflexivity to uncover the deeply embedded values, incentives, and sociotechnical imaginaries that shape the field of IR. It recognizes the agency of the field and its practitioners in shaping the world, while respecting that the challenges and necessary interventions are both fundamentally sociotechnical in nature, and therefore the theories and practices of this field must be co-developed with cross-disciplinary collaboration with social science and humanities scholars, legal and policy experts, civil rights and social justice activists, and artists, among others.
Emancipation research in other fields.
To develop an emancipatory IR research agenda, we can take inspiration from other scholarly fields that have incorporated humanistic values and emancipatory aspirations. Young et al. [144] surveyed the body of emancipation research in information science, and identified four components of emancipation prevalent in the literature: (i) Agency (i.e., freedom to act), (ii) dialogue (i.e., freedom to express), (iii) inclusion (i.e., freedom to belong), and (iv) rationality (i.e., freedom to think). Wright [140] defines emancipatory social science as an intellectual enterprise seeking to “generate scientific knowledge relevant to the collective project of challenging various forms of human oppression”. Wright outlines a framework that enumerates three specific tasks for the field: (i) Systematic diagnosis and critique of the world as it exists, (ii) envisioning viable alternatives, and (iii) elaborating a theory of social transformation. According to Wright, emancipatory social science is “a theory of a journey from the present to a possible future: the diagnosis and critique of society tells us why we want to leave the world in which we live; the theory of alternatives tells us where we want to go; and the theory of transformation tells us how to get from here to there – how to make viable alternatives, achievable”.
In human–computer interaction (HCI), Bardzell and Bardzell [11, 10] define humanistic HCI as research and practices in the field that deploy humanistic epistemologies and methodologies. They note that emanicipatory HCI is “a fundamental goal of virtually all humanistic HCI”. Related, several strands of HCI research incorporate anti-oppressive and emancipatory [115], feminist [12, 14, 16, 15, 13], queer [81, 77], postcolonial and decolonial [67, 43, 119, 6, 68], anti-racist [4], anti-casteist [124, 125], anti-ableist [137, 118], post-capitalistic [48, 24], and anarchist [74, 82] epistemologies, theories and practices.
Similarly, in the fields of machine learning (ML) and artificial intelligence (AI) there are several works that call for anti-oppressive and emancipatory [72, 110], anti-fascist [88], decolonial [5, 93], anti-casteist [71, 109], and abolitionist [19, 9, 45, 138] perspectives and design interventions. In data science, D’ignazio and Klein [42] and Guyan [57] discuss how feminist and queer epistemologies can inform the field, and Monroe-White [95] adopted Wright’s framework [140] to propose an emancipatory agenda for data science to mitigate the harms to marginalized populations.
3 Practices, projects, and provocations
To develop an emancipatory research agenda for IR, we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. However, the list of practices, projects, and provocations discussed here are neither immutable nor complete. It is ultimately the responsibility of the community engaged in emancipatory research to regularly reflect on, produce evidence-based critique, challenge, and reshape its agenda to affect tangible progress in our collective struggles against dehumanization, oppression, and marginalization. It may be worth noting at this stage that due to the saliently sociotechnical nature of the challenges in this area, the practices, projects, and provocations too will rarely reside strictly in the domain of computing and technology and must be informed by inter-disciplinary perspectives and lived experiences of marginalized peoples. We illustrate the practices, projects, and provocations described in this section in Figure 1.
3.1 Practices of emancipatory IR
We adapt Wright’s emancipatory social science framework [140] to develop a framework for the practices of emancipatory IR. Like Wright’s, our framework consists of diagnosing and critiquing, imagining viable alternative futures, and elaborating our theories of change.
Diagnose and critique.
The starting point for emancipatory IR research is identifying the ways in which existing IR methods and systems, the cost of IR research and development, and the arrangements within the IR community may contribute systemic harms to peoples, act contrary to aspirations of social and political justice, and impede emancipatory struggles of the peoples. These harms may include:
  • Suppressing voices of the oppressed and marginalized
  • Enabling surveillance and public opinion manipulation
  • Imposing representational harms through derogatory portrayal, stereotyping, or erasure in presentation of retrieved results
  • Imposing allocative harms to groups through under-exposure of their content in retrieved results or under-exposing groups to socioeconomic opportunities
  • Exploitation of ecological resources and labor for IR research and development that mirror racial capitalism and coloniality
  • Capture of IR research by the military–industrial complex that pushes the field to disproportionately prioritize economic and military interests over concerns of knowledge production, public health education, and information literacy, and produce technologies that serve as tools of oppression (e.g., for surveillance)
  • Concentration of extreme power and wealth among few individuals and institutions in ownership of popular information access platforms
  • Lack of representation and diversity within the IR community leading to propping up colonial cisheteropatriarchal capitalist values and erasure of feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist perspectives
Imagine viable alternative futures.
The second practice of emancipatory IR is to imagine and develop desirable, viable, and achievable alternatives to our IR technologies and ways of organizing our communities with the aim of addressing the harms and injustices identified in the diagnosis and critique stage. Mitra [91] argues that the IR community should explicitly articulate the sociotechnical imaginaries [69, 70] that influence the design and development of IR technologies and platforms. He enquires “what would IR systems look like if designed for futures informed by feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts, and if the focus of IR research was not to prop up colonial cisheteropatriarchal capitalist structures but to dismantle them?” This is a call for us to imagine post-oppressive worlds, and the role of IR both in that world and in the journey that leads us there.
Elaborate theories of change.
Our final practice concerns with the development of theories of change [133, 23, 122, 134] that makes explicit how we aim to realize the alternative futures in the face of current realities and how our own work contributes towards those goals. This should be grounded in theories of power and co-developed with cross-disciplinary scholars, legal and policy experts, activists, artists, and others to realize real structural changes. This involves both developing new research agendas as well as experimenting with and realizing new arrangements within the IR community and critically reflecting on our relationships with other institutions (e.g., industry and government). To provide concrete examples, we next discuss potential projects that we believe emancipatory IR should invest in.
Figure 1: Illustrating the practices, projects, and provocations of emancipatory IR. The hexagons at the center represent several potential projects. The empty hexagons indicate that the set of mentioned projects are not complete, and the community should over time identify new ones and shape existing ones, as appropriate. Surrounding the projects is the framework of practices of emancipatory IR that are relevant to all projects in this area, consisting of (i) diagnose and critique, (ii) imagine viable alternative futures, and (iii) elaborate theories of change. Provocations to spark news ideas and research directions are depicted as a nudging hand.
3.2 Emancipatory projects in IR
Next, we enumerate some of the potential projects relevant to emancipatory IR. Such examples are useful for readers to get more concrete sense of the aspirations and type of work relevant to this field, and can also serve as a starting point for developing a more thorough research agenda. Ultimately, the direction and what specific project we invest in must be decided and iterated on over time by those who see themselves as part of this community, which we must ensure includes more than just IR researchers and technologists as we emphasized previously.
Safeguard against capture.
A central concern of emancipatory IR should be to ensure that our information access platforms cannot be easily captured by few privileged individuals and institutions or authoritarian regimes, and also to ensure that access to information services cannot easily be blocked by authoritarian regimes in times of protests to suppress dissent. These concerns and initiatives are timely in the face of the 2024 edition of the World Economic Forum’s Global Risks Report [2] that calls out “technological power concentration” as one of the top global risks for the upcoming decade, and broader growing concerns about the alignment of tech industry with authoritarianism and fascism [73, 51, 126, 78, 44]. In the social media landscape, there has been recent announcements [86, 3] of formation of new non-profit foundations for managing key social media ecosystem and platforms, specifically with the goal to safeguard against billionaires’ control over our digital public squares. In IR, there has also been recent investments to create open infrastructures for internet search, such as the Open Web Search project. [^4] https://openwebsearch.eu/
While decentralization of IR infrastructure and federation may be critical for protecting against capture, it is also important to acknowledge that in a world shaped by white supremacy, colonialism, patriarchy, cisnormativity, heteronormativity, neuronormativity, and casteism as organizing political forces, the decentralization of platforms in themselves do not constitute sufficient conditions for creating safe and welcoming online spaces for historically marginalized peoples nor to ensure these platforms align with our emancipatory aspirations [64]. Instead, the challenges that we face requires both decentralized infrastructure and technology innovations, as well as experimentation with and formation of new global organizational structures for the ownership, management, and moderation of these platforms to ensure that the voices of the oppressed and marginalized are uplifted, not suppressed.
Safeguard against surveillance.
IR applications—such as web search, recommender systems, and targeted digital advertising—have been major drivers of surveillance capitalism [146]. An important commitment IR can make towards anti-oppression is to ensure that the field is neither creating the technologies and infrastructures for surveillance, nor creating additional demands for such infrastructures out of commercial interests. User behavior data have been the “secret” sauce for many successful commercial IR platforms. For example, in search, user interactions with search result pages are continuously logged and modeled at incredible scale to provide better search relevance to customers. It is therefore not surprising that modeling of user clicks [33] and even cursor movements [40] have historically garnered lots of interest in IR research. However, this creates demands for unscrupulous surveillance of users on these platforms. In response, different regulatory safeguards have been proposed, including notably the principles of data minimization defined in Article 5 of the European Union’s General Data Protection Regulation (GDPR)  [107].
To address this, the IR community must imagine alternative ways of building its systems that does not necessitate such large scale data collection to simply make these platforms usable. We must both push the boundaries of effective retrieval methods in the absence of user behavior data and develop new protocols for collection, storage, and application of user data that implements strict safeguards for user privacy and require stricter and meaningful user consent for access by the platforms. We need to explore both technological innovations (such as, homomorphic encryption) as well policy safeguards and changing social norms around data stewardship and transparency of how institutions collect, store, and use data, and implement concrete safeguards to ensure that authoritarian governments and their institutions cannot get access to this data which may enable harassment and intimidation of activists, journalists, and marginalized populations. Lastly, we should be cognizant of potential dual-use of IR technologies—such as the impact of improved cross-lingual search on surveillance of foreign individuals, institutions, and governments—and be critical of funding for such research that comes from governments and security agencies.
Safeguard against manipulation.
Misinformation and disinformation pose serious threats to democracies worldwide [80, 46, 46, 79]. The 2024 edition of the World Economic Forum’s Global Risks Report [2] ranks misinformation and disinformation as the top global risk during the subsequent two years. A study spanning six years, 26 countries, and several election periods find that political misinformation is particularly salient in current wave of radical right populism and its opposition to liberal democratic values and institutions [123]. Curtailing spread of online misinformation and disinformation that aims to dehumanize immigrants, the poor, trans and nonbinary, and racially marginalized peoples, is critical to our social justice efforts and our effort to build solidarity with each other.
Beyond misinformation, we must also pay close attention to new vectors of public manipulation that may become plausible in near future based on our current technological trends. A particularly worrying development in this area is the recent emergence of “persuasive AI” [27, 29, 101, 47]. Mitra et al. [92] warn that the massive trove of detailed data on user behavior and preferences combined with the capabilities of generative AI to produce persuasive language and visualizations could create tools of mass manipulation and poses serious risks to functioning of global democracies. Emancipatory IR researchers must engage in evidence-based critique of such technology development and push for policies and community norms that push back against research in support of such applications within the field before such applications of AI materialize.
Safeguard against dehumanization and inequitable outcomes.
Emancipatory IR should concern itself with ensuring that IR systems do not lead to representational harms (e.g., denigration, stereotyping, and erasure) of historically marginalized peoples and inequitable outcomes which may further reinscribe historical sociopolitical and economic oppression and exploitation. Over the last several years, there has been several works on fairness in IR. These have typically focused on fairness of quality-of-service (e.g.[89, 90, 97, 142]) and fairness of exposure (e.g.[20, 113, 114, 41, 145, 103, 141]). However, there is less clarity on how much this body of fairness research has practically influenced material changes in popular IR platforms [91]. Fairness research in the emancipatory context must move beyond just formalization and theorization of fairness; it must recognize and be grounded in the understanding of historical injustices and structural violence, and must ensure that this research opposes dehumanization and material inequities in the real world. This includes also contending with the politics of classification [38] and critiquing practices in fairness research, such as in context of gender fairness [105], that may themselves perpetuate further harm.
We must also hold institutions that own and operate IR platforms accountable to measurable equitable and humanistic outcomes. If an institution invests in fairness research but do not operationalize them in their products, then that research only serves to launder institutional reputation and is counter to our real emancipatory goals. To hold institutions accountable, our research agenda should also include challenges of auditing IR platforms and also explore other regulatory, sociopolitical, and technological mechanisms of accountability that is grounded in the recognition of the structural power inequities that make it challenging to do so.
Promote emancipatory pedagogy.
Emancipatory IR must abandon any false notions of neutrality between the oppressor and the oppressed, and develop frameworks situated in theories of power to uplift marginalized voices and make spaces for emancipatory pedagogy. It must expressively refuse to serve as a soapbox for the powerful and wealthy. To make this further challenging, we must do so without concentrating the power to define oppressor-oppressed relations in the hands of platform owners. Can we, for example, learn from Freire’s [50] anti-oppressive pedagogy to imagine IR systems that do not merely retrieve but provide spaces for dialogical interactions between information seekers? Can we build communities of experts who can provide context from critical scholarship to inform search results and aids in our collective sense-making and solidarity-building on topics relevant to our emancipatory struggles? What if we go beyond fixing under-representation of marginalized groups in, say, image search results for occupational roles and reclaim those digital spaces as sites of resistance and emancipatory pedagogy [91]?
Emancipatory IR must also critically interrogate the relationship between information access platforms and the publishers and news media. Kim et al. [75] find media indoctrination as one of the strongest predictors of autocratic survival. In light of these dynamics, we must reexamine how platforms can support local newsrooms that often do critical work in reporting on issues relevant to marginalized populations. In this context, support must include both ensuring these publishers get exposure in search results as well as innovations in platform business models that can support local newsrooms financially, such as through revenue sharing.
Visibilize and mitigate ecological and human costs of IR research and development.
We must also be concerned of the cost of doing IR research and operating IR platforms. This includes ecological costs [111, 147], such as energy and water consumption and harmful emissions, as well as the appropriation of data labor and impact on the wellbeing of workers. Belkhir and Elmeligi [18] estimate that Information and Communications Technology industry on the whole will account for 14% of global emissions by 2040. In the US, data centers are projected to consume around 6% of the total national electricity by 2026 [59]. Appropriation of data labor includes both uncompensated appropriation of works by writer, artists, and programmers [34, 35, 129, 128, 112, 25, 26, 131, 130, 8, 85, 32, 31, 53, 132] and under-paid crowdwork for data labeling [104, 135, 121, 7, 61, 143, 62]. The latter mirrors historical patterns of global labor exploitation and racial colonial capitalism [60, 21, 99, 76, 36, 96, 120], and have been associated with severe mental health cases among crowdworkers [22, 62]. Other concerns include land use for data centers and how that affects communities living on those lands and indigenous sovereignty [108, 127].
Emancipatory IR must invest in accurate accounting, reporting, and reduction of the ecological and human cost of our research. This includes the cost of training models and experimentation, the cost of human annotations, and the cost of air travel for conference attendance, among others. It includes a commitment towards not pursuing research directions whose cost include accelerating climate change and broader adoption of virtual conferencing to minimize air travel. It also includes the recognition that these costs cannot just be minimized by building more efficient models [^5] For example, because of Jevons paradox: https://en.wikipedia.org/wiki/Jevons_paradox. or automated labeling methods, and that the solution as always is at the intersection of the social, the political, the economic, and the technological.
Community building and organizing.
The work of building an emancipatory research agenda must start with building an emancipatory research community. We need to create safe spaces where researchers from the IR community can engage and meaningfully interact with social and political science scholars, critical theorists, legal and policy experts, activists, and artists to co-develop an agenda for this field. We must reflect on the current social and organizational arrangements within the field and our relationships with other institutions, such as industry and government. We must take on the difficult challenge of critiquing the funding mechanisms that enable IR research and its role in privileging certain areas of research and technology development that are antagonistic to our emancipatory aspirations and social justice. We must reimagine these arrangements, build counter-structures, and establish new means of funding emancipatory IR research that can sustain itself in the face of attack from the privileged, powerful, and wealthy. We must nurture a culture in our community that recognizes direct actions that challenge power, such as demonstrating and whistleblowing [65], and create support mechanisms to protect those individuals. We must ensure that our community is inclusive and representative of marginalized peoples and that we view their lived experiences with structural oppression as critical to informing research in our field. And, we should view emancipatory IR as situated in broader movement building against oppression and marginalization.
3.3 Provocations
Emancipatory IR must leverage instruments of provocation to challenge our thinking and help us reimagine radically alternative futures. The inception of emancipatory IR must not be in a vacuum, but must be informed by emancipatory ideologies borne of our historical and ongoing struggles, including feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts. Our challenge is not only to imagine radically new emancipatory futures, but also practicing those desired values and build prefigurative counter-structures both in the form of new technological designs and sociopolitical rearrangements within our communities. The task to change technology is inextricably linked to the task of changing ourselves and our own social organizations. The practice of reflexivity becomes important in this context to articulate (and critique if necessary) our own values, positionality, and incentives, as well as our relationships with institutions and social structures that we in turn intend to change. It is in this process that we desperately need to engage scholars outside of computing, learn from generations of activists who have been engaged in emancipatory struggles, and let ourselves be influenced by the artists and writers that dare to imagine radically free futures and humanistic social arrangements.
4 Discussion
4.1 The emancipatory in emancipatory IR
The term “emancipation” originally described the abolitionist struggles against slavery, but later came to signify a broader vision of dismantling all forms of oppression. Emancipatory thinking in humanities have manifested in many forms, including Marxism, Critical Theory, feminism, and post-colonialism. Emancipatory IR aims to be informed by and incorporate these different theories and epistemic frameworks in the conception and design of information access methods and systems. For emancipatory IR to be authentically emancipatory, this work must not be simply limited to an intellectual exercise. Our scholarship must be piece of a broader movement building grounded in ongoing struggles in our world against dehumanization, marginalization, and oppression. Our critiques must be followed by push for dismantling oppressive structures and building counter-structures in its place. Our theories must be put into practice and tested in the real world, and our success measured in concrete emancipatory outcomes. Access to information is critical to our collective sense-making of our place in this world and our relations with it, and it is exactly why throughout history, from print media to present-day social networks, authoritarians forces have tried to control what and how information is disseminated. Emancipatory IR stands in solidarity with the oppressed and their struggles, and accepts the challenge of building the information infrastructure for resistance.
4.2 The IR in emancipatory IR
But what makes emancipatory IR truly IR? How are questions of emancipation relevant to this field of computing whose research typically concerns with questions of ranking, evaluation, efficient system design, language modeling, human information interaction, and other similar topics? To this we respond by challenging the reader to reflect on how each of these important concerns in IR intersect with the examplar projects we discussed in Section 3.2. What new research questions emerge in context of indexing and ranking when our IR platforms are federated and decentralized? What new research challenges emerge in human information interaction if our goal is to promote emancipatory pedagogy? How do we mitigate the harms from appropriation of data labor that is so very prevalent in our modern-day language modeling approaches? How do we reimagine efficiency and evaluation when our aspirations go beyond fetching information artifacts at lightning speed to encouraging deeper collective understanding and new knowledge production that are key to our social progress? With these questions our aim is to challenge the field to abandon a more restrictive view of what IR research is which may limit itself to algorithmic advances as measurable on shared benchmarks, and adopt a more expansive view of IR that aims to situate information access as a vehicle for social good.
5 Conclusion
This piece is a provocation for the IR community to situate itself in the generational emancipatory struggles and stand in resistance to the growing epidemic of authoritarianism and fascism that we are experiencing worldwide today. While focused on IR, it is a challenge to the broader computing and technology community to reimagine our work to not serve the privileged few by further concentrating immense power and wealth, but as a liberating force that connects us and humanizes us universally. Technology after all is a project of world-making. So, what kind of a world are we collectively going to build together?